OpenStudy (anonymous):

fifity two of 80 marbles are red. what percent if marbles are red

OpenStudy (mathhelp346):

65

OpenStudy (anonymous):

52/80 = .065 then multiply the result by 100

OpenStudy (mathhelp346):

52/80= 26/40=13/20. multiple both by 5 and you get 65%